A. How many millimoles of magnesium chloride are there in 4 liters of 0.2 M MgCl2

a. How many nanomoles are in 2.3 x 10-1 μmoles?
Answer #1

A)

4 liters of 0.2 M MgCl2 = volume of solution in L * molarity = 4.0 L * 0.2 mole / L = 0.8 mole.

we know,

1 mole = 1000 mmole.

thus

0.8 mole = 0.8 * 1000 = 800 mmole (answer)

a)

we know,

1 μmole = 1000 nanomoles

thus

2.3 x 10-1 μmoles =  2.3 x 10-1 * 1000 = 230 nanomoles (answer)